Find the equation of the line through (2,3) and parallel to 3x-2y=7?

Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

3 x + -2 y = 7
Find the slope of this line
make y the subject
-2 y = -3 x + 7
Divide by -2
y = 1.5 x -3.5
Compare this equation with y=mx+b
slope m = 1.5 1 1/2
The slope of a line parallel to the above line will be the same
The slope of the required line will be 1.5
m= 1.5 ,point ( 2 , 3 )
Find b by plugging the values of m & the point in
y=mx+b
3 = 3.00 + b
b= 0 0
m= 1.5 1 1/2
Plug value of the slope and b in y = mx +b
The required equation is y = 1.5 x
(y = 3/ 2 )
m.ananth@hotmail.ca

Answer by stanbon(75887)

Find the equation of the line through (2,3) and parallel to 3x-2y=7?
---------
Find the slope of the given equation:
3x-2y = 7
2y = 3x-7
y = (3/2)x - (7/2)
----
slope = 3/2
-----
Any line parallel to it must have slope = 3/2
-----
Form: y = mx + b
m = 3/2 and y = 3 when x = 2
Solve for "b":
3 = (3/2)*2 + b
b = 0
----
Equation:
y = (3/2)x
=================
cheers,
Stan H.
=================